OpenStudy (anonymous):

what is the principal value of cos^-1{cos 23pi/20}

OpenStudy (anonymous):

\[\frac{17\pi}{20}\]

OpenStudy (anonymous):

i think, let me check with a calculator

OpenStudy (anonymous):

yeh its correct

OpenStudy (anonymous):

its an interesting method to solve , Ill do a quick picture now on paint and hopefully try to explain

OpenStudy (anonymous):

ok..m waitin

OpenStudy (anonymous):

Well, not the best picture , but good enough . Now, first thing to notice is that the answer is not 23pi / 20 , because 23pi/20 is not in the range of cos^-1

OpenStudy (anonymous):

we must get the angle into the interval 0<x<pi

OpenStudy (anonymous):

and 23pi/20 is a bit more than pi

OpenStudy (anonymous):

hey can you tell me what is meant by principal value?? i dont know if i get the idea u knw

OpenStudy (anonymous):

however, the cos graph is symmetrical about the minimum points , that is cos(23pi /20) does equal the same value as cos (17pi /20) because they are both the same distance from the minimums

OpenStudy (anonymous):

principal value just means , "your answer must be in the range 0<x<pi" pretty much only then does cos^-1(cos(x)) = x

OpenStudy (anonymous):

if you really want to check your understanding then try \[\sin^{-1} (\sin(\frac{23\pi}{20}))\]

OpenStudy (anonymous):

its a very similar idea

OpenStudy (anonymous):

using translations of the graph , you get the angle into the range -pi/2 to pi/2 ( because the range of sin^-1 is -pi/2 to pi/2 )

OpenStudy (anonymous):

then apply sin^-1(sin(x)) = x

OpenStudy (anonymous):

hmm.. i understand what you are saying. i would like to know how you got the value 17pi/20? i know its at a distance pi/4 from 23pi/4. i wanna know if there is some other way to work it out.

OpenStudy (anonymous):

there is other ways of working it our but I dont remember them at this point in time

OpenStudy (anonymous):

hey is the prin. value for \sin^{-1} (\sin(\frac{23\pi}{20})) 13 pi/20?

OpenStudy (anonymous):

\[\sin^{-1} (\sin(\frac{23\pi}{20})) \]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

draw a picture of a sin graph this time

OpenStudy (anonymous):

yea i did..

OpenStudy (anonymous):

my picture is not good, but you get the idea

OpenStudy (anonymous):

23pi / 20 , is a bit more than pi , yes?

OpenStudy (anonymous):

now, we need to get the angle 23pi/20 into the range -pi/2 <= <= pi/2

OpenStudy (anonymous):

okk...

OpenStudy (anonymous):

isnt it at a distance pi/2 from it?

OpenStudy (anonymous):

now, I know the graph of sin is symmetric about pi/2

OpenStudy (anonymous):

so if I find the distance between 23pi/20 and pi/2 , then this is the distance I must go to the left and I will still have the equivalent value of the function

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

now, the distance between pi/2 and 23pi/ 20 = (23/20 - 1/2) pi = 13pi/20

OpenStudy (anonymous):

this means that if I go 13pi/20 units left from pi/2 , then I would be at a point that is a reflection in the line x=pi/2

OpenStudy (anonymous):

ie sin(23pi / 20 ) = sin ( pi/2 - 13pi/20 ) = sin ( -3pi/20 )

OpenStudy (anonymous):

now, since the angle of the sin function is in the range of the invese sin function

OpenStudy (anonymous):

I can apply sin^-1(sin(x) ) = x

OpenStudy (anonymous):

so our answer is -3pi/20

OpenStudy (anonymous):

oh..i understand now. thankyou so much=)

OpenStudy (anonymous):

which can be checked on the calculator , put cal in radian mode , type in " sin^-1(sin(23pi /20) )"

OpenStudy (anonymous):

with the brackets, and you will get something like -0.47 .... which is equal to -3pi/20

OpenStudy (anonymous):

ya..got it